OpenStudy (anonymous):

Find the equation of the quadratic function with zeros -1 and 1 and a vertex at (0, -4). is it y = x2 − 1?

OpenStudy (amistre64):

we start by filling in the basic form (x+1)(x-1) = x^2 -1 Now we plug in x = 0 from your point to get; 0^2 -1 = -1 ; this is the basic parabola at 0, we need to multiply this to get to -4 -1n = -4 n = 4; now multiply the whole thing by 4 to get the answer 4(x^2 -1) = 4x^2 -4